OpenStudy (shootingstar28):

A rectangular window has a length of 12 ft and a width of x ft. Which expression represents the area of the window? 12/x 6x 12x x + 12

OpenStudy (rayep12):

c

jimthompson5910 (jim_thompson5910):

Agreed. The area is equal to the length (12) times the width (x) area = length times width = L*W = 12*x = 12x